r2 - 24 May 2011 - 13:40:16 - KaushikDeYou are here: TWiki >  Admins Web > MinutesDataManageMay24



Minutes of the US ATLAS Data Management meeting, May 24, 2011
  • Previous meetings and background : IntegrationProgram
  • Coordinates: Tuesdays, Noon Central
    • (570) 310-0130, Access code: 735188; Dial *6 to mute/un-mute.


  • Meeting attendees: Saul, Wei, Hiro, Armen, Charles, Wensheng, Shawn, Bob, Dave, Patrick,
  • Apologies:
  • Guests:

Topics for this week

  • Storage status - Armen
    • Everything looks good.
    • IU is being migrated. Within a month or two the site and tokens will disappear.
    • In the US, sites will continue to do PRODDISK cleanup (_dis datasets).
    • SWT2 is now reporting space per token using Bestman 2.
  • Central deletion - Hiro, Armen, Charles
    • Rate has been low at BNL? Currently only 1-2 Hz. Still under discussion with central deletion.
    • Need to cleanup dark data on MCDISK. Wait till migration is completed.
  • Storage reporting - Hiro
    • unallocated, unpowered tokens.
      • SWT2 - done, MWT2 - next week, NET2 - wait for Bestman 2, BNL - done?
      • Hiro will add these tokens to his monitoring - next week?
  • Hot issues
  • AOB
    • SWT2 GROUPDISK 133TB used - more than twice other sites? Ok - busy tokens.
    • Patrick - Bestman 2 at SWT2_CPB. Completed.
    • Hiro - deployment of new SS at BNL. Completed.
    • Tier 3 consistency reports - who and how fixes them?

-- KaushikDe - 24 May 2011

About This Site

Please note that this site is a content mirror of the BNL US ATLAS TWiki. To edit the content of this page, click the Edit this page button at the top of the page and log in with your US ATLAS computing account name and password.


Powered by TWiki
This site is powered by the TWiki collaboration platformCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback